### v1

@@ Abstract
Deep neural networks for medical image segmentation often produce overconfident results misaligned with empirical observations. Such miscalibration, challenges their clinical translation. We propose to use marginal L1 average calibration error (mL1-ACE) as a novel auxiliary loss function to improve pixel-wise calibration without compromising segmentation quality. We show that this loss, despite using hard binning, is directly differentiable, bypassing the need for approximate but differentiable surrogate or soft binning approaches. Our work also introduces the concept of dataset reliability histograms which generalises standard reliability diagrams for refined visual assessment of calibration in semantic segmentation aggregated at the dataset level. Using mL1-ACE, we reduce average and maximum calibration error by 45% and 55% respectively, maintaining a Dice score of 87% on the BraTS 2021 dataset. We share our code here: https://github.com/cai4cai/ACE-DLIRIS.
@@ 1 Introduction
Deep neural networks (DNNs) have significantly advanced the field of semantic segmentation. However, their large capacity, makes them susceptible to overfitting, leading to overconfident predictions that do not accurately reflect the underlying uncertainties inherent to the task [1]. In medical image segmentation, where the confidence of predictions can be as crucial as the predictions themselves, and overfitting is exacerbated by smaller datasets, such overconfidence poses significant risks, particularly when this miscalibration is not communicated to the end-user. Appropriate management of confidence can also have benefits for patient outcomes, for example in radiotherapy contouring [2].
The Dice Similarity Coefficient (DSC) loss, is a popular choice in medical image segmentation due to its robustness against class imbalance [3], but it is known to yield poorly calibrated, overly confident predictions [4].
@@
Traditional post-hoc calibration strategies, such as temperature scaling (Ts) [1, 7], offer a partial remedy by adjusting confidence levels uniformly. While more nuanced post-hoc calibration methods offer class-specific adjustments [8, 9, 10], they still fail to exploit the extensive parameter space of DNNs [11, 12].
@@
In this work, we present a novel auxiliary loss function, marginal L1 Average Calibration Error (mL1-ACE), designed to enhance calibration accuracy in medical image segmentation without compromising on segmentation quality. This represents the first clear attempt to use established calibration metrics directly as an auxiliary loss in the context of medical image segmentation. Despite presumptions related in previous DCA-based methods, our approach is inherently differentiable, even with hard-binning of probabilities, eliminating the need for surrogates or soft-binning techniques. mL1-ACE employs a class-wise calibration strategy, opting for ACE over ECE to better address the class imbalances in medical datasets. When combined with standard loss functions, in particular DSC, mL1-ACE demonstrates superior calibration performance on the BraTS2021 dataset [17], significantly surpassing conventional post-hoc calibration methods like Ts. Additionally, we introduce dataset reliability histograms, offering a comprehensive visual tool for evaluating and communicating model calibration across image segmentation datasets, enhancing transparency in reporting model reliability beyond predictive performance.
@@ 2 Method
where $n^{c}_{m}$ is the number of in bin $B^{c}_{m}$, the $m^{\textrm{th}}$ bin for class $c$, and $N$ is the total number of voxels for an image. More formally, we define $o^{c}_{m}$ and $e^{c}_{m}$ as:
@@ 4 Results
Figure 3: Comparative analysis of DSC against calibration metrics for average tumour component of BraTS 2021 dataset. Statistically significant difference (p < 0.01).
@@
Loss Function
ET
TC
WT
Avg
CE
0.82 ± 0.19
0.87 ± 0.18
0.90 ± 0.11
0.86 ± 0.14
CE + L1-ACE
0.80 ± 0.20
0.86 ± 0.19
0.87 ± 0.12
0.84 ± 0.15
DSC
0.84 ± 0.17
0.87 ± 0.20
0.89 ± 0.12
0.87 ± 0.14
DSC + L1-ACE
0.83 ± 0.18
0.88 ± 0.18
0.89 ± 0.11
0.87 ± 0.13
DSC-CE
0.85 ± 0.17 0.89 ± 0.17 0.90 ± 0.11 0.88 ± 0.13
DSC-CE + L1-ACE
0.82 ± 0.20
0.86 ± 0.21
0.88 ± 0.13
0.85 ± 0.16
We show the effect of Ts and mL1-ACE on the calibration of the models in Table 2. Our auxiliary loss has little effect on the calibration of the model trained with CE. Our loss shows the biggest improvement in calibration when paired with the DSC loss function, with a 45% reduction in ACE (p < 0.01). Overall the use of Ts shows marginal improvement in calibration, with the exception of CE loss where it has a small detrimental effect. Similar conclusion can be drawn for MCE and ECE metrics, shown in Table A1 and Table A2 in the appendix. We also note that the number of bins used to calculate our calibration metrics, has little effect, with ACE increasing by less than 3% when increasing the bins from 10 to 100 (p < 0.01), as shown in Figure A1 in the appendix.
@@ Acknowledgments
Appendix
Table A1: Maximum Calibration Error
ET
TC
WT
Avg
CE
0.26 ± 0.19
0.28 ± 0.18
0.23 ± 0.12
0.25 ± 0.12
CE + Ts
0.30 ± 0.16
0.29 ± 0.16
0.25 ± 0.10
0.28 ± 0.10
CE + L1-ACE
0.24 ± 0.19
0.26 ± 0.18
0.20 ± 0.15
0.23 ± 0.14
CE + L1-ACE + Ts
0.23 ± 0.15 0.25 ± 0.16
0.20 ± 0.12
0.23 ± 0.11
Dice
0.59 ± 0.15
0.61 ± 0.15
0.55 ± 0.13
0.58 ± 0.11
Dice + Ts
0.56 ± 0.15
0.58 ± 0.15
0.51 ± 0.14
0.55 ± 0.11
Dice + L1-ACE
0.29 ± 0.18
0.30 ± 0.17
0.23 ± 0.14
0.27 ± 0.13
Dice + L1-ACE + Ts
0.23 ± 0.17 0.25 ± 0.16
0.21 ± 0.11
0.23 ± 0.12
Dice-CE
0.46 ± 0.17
0.46 ± 0.17
0.36 ± 0.17
0.43 ± 0.13
Dice-CE + Ts
0.26 ± 0.19
0.32 ± 0.19
0.19 ± 0.12
0.26 ± 0.13
Dice-CE + L1-ACE
0.29 ± 0.21
0.30 ± 0.20
0.21 ± 0.14
0.27 ± 0.15
Dice-CE + L1-ACE + Ts
0.25 ± 0.19
0.27 ± 0.18
0.20 ± 0.11
0.24 ± 0.13
Table A2: Expected Calibration Error (×$\times 10^{-4}$)
ET
TC
WT
Avg
CE
4.21 ± 7.49
5.66 ± 10.5
9.71 ± 11.3
6.53 ± 7.53
CE + Ts
7.77 ± 7.16
10.2 ± 10.0
19.1 ± 9.45
12.4 ± 6.93
CE + L1-ACE
4.00 ± 4.75
6.16 ± 9.41
12.0 ± 12.9
7.38 ± 6.94
CE + L1-ACE + Ts
21.1 ± 4.69
27.2 ± 6.65
62.0 ± 11.1
36.8 ± 5.69
Dice
5.09 ± 7.67
7.48 ± 16.5
15.5 ± 17.8
9.34 ± 11.1
Dice + Ts
4.95 ± 7.63
7.36 ± 16.5
15.0 ± 17.8
9.10 ± 11.1
Dice + L1-ACE
3.47 ± 7.56
4.50 ± 8.45
10.1 ± 12.8
6.03 ± 6.84
Dice + L1-ACE + Ts
3.27 ± 7.17 4.41 ± 8.00
12.0 ± 10.9
6.54 ± 5.84
Dice-CE
4.40 ± 7.57
5.70 ± 12.2
11.6 ± 14.4
7.21 ± 8.62
Dice-CE + Ts
3.84 ± 7.43
5.20 ± 11.8
10.2 ± 13.4
6.41 ± 8.11
Dice-CE + L1-ACE
3.91 ± 8.09
6.04 ± 14.6
10.3 ± 14.8
6.75 ± 9.84
Dice-CE + L1-ACE + Ts
4.23 ± 7.68
6.56 ± 14.0
13.9 ± 12.6
8.23 ± 8.83
@@
Figure A1: Comparison of calibration metrics using different number of bins when evaluated on the DSC + mL1-ACE loss model. Both ACE and ECE are very stable with respect to number of bins (5, 10, 20, 50, 100, 1000). MCE shows a larger increase from 100 to 1000 bins, this is due to the presence of empty bins, a likely occurrence when using such a high number of bins.

### v4

@@ Abstract
Deep neural networks for medical image segmentation often produce overconfident results misaligned with empirical observations. Such miscalibration challenges their clinical translation. We propose to use marginal L1 average calibration error (mL1-ACE) as a novel auxiliary loss function to improve pixel-wise calibration without compromising segmentation quality. We show that this loss, despite using hard binning, is directly differentiable, bypassing the need for approximate but differentiable surrogate or soft binning approaches. Our work also introduces the concept of dataset reliability histograms which generalises standard reliability diagrams for refined visual assessment of calibration in semantic segmentation aggregated at the dataset level. Using mL1-ACE, we reduce average and maximum calibration error by 45% and 55% respectively, maintaining a Dice score of 87% on the BraTS 2021 dataset. We share our code here: https://github.com/cai4cai/ACE-DLIRIS.
@@ 1 Introduction
Deep neural networks (DNNs) have significantly advanced the field of semantic segmentation. However, their large capacity makes them susceptible to overfitting, leading to overconfident predictions that do not accurately reflect the underlying uncertainties inherent to the task [8]. In medical image segmentation, where the confidence of predictions can be as crucial as the predictions themselves, and overfitting is exacerbated by smaller datasets, such overconfidence poses significant risks, particularly when this miscalibration is not communicated to the end-user. Appropriate management of confidence can also have benefits for patient outcomes, for example in radiotherapy contouring [2].
The Dice Similarity Coefficient (DSC) loss is a popular choice in medical image segmentation due to its robustness against class imbalance [15], but it is known to yield poorly calibrated, overly confident predictions [16].
@@
Traditional post-hoc calibration strategies such as temperature scaling (Ts) [8, 21] offer a partial remedy by adjusting confidence levels uniformly. While more nuanced post-hoc calibration methods offer class-specific adjustments [13, 6, 11], they still fail to exploit the extensive parameter space of DNNs [14, 9].
@@
In this work, we present a novel auxiliary loss function, marginal L1 Average Calibration Error (mL1-ACE), designed to enhance calibration accuracy in medical image segmentation without compromising on segmentation quality. This represents the first clear attempt to use established calibration metrics directly as an auxiliary loss in the context of medical image segmentation. Despite presumptions related in previous DCA-based methods, our approach is inherently differentiable, even with hard-binning of probabilities, eliminating the need for surrogates or soft-binning techniques. mL1-ACE employs a class-wise calibration strategy, opting for ACE over ECE to better address the class imbalances in medical datasets.
While the concept of average calibration error (ACE) is not new, our key contribution lies in adapting it as a directly differentiable auxiliary loss for semantic segmentation, a novel approach not addressed in prior literature. This provides a major distinction from the use of ACE as a metric in classification tasks. The transition from classification to segmentation is a key aspect which we are the first to identify, enabling ACE to be used as a loss rather than a metric. Classification training pipelines provide too few samples within a mini-batch to meaningfully compute its ACE, in contrast to segmentation where each voxel provides a sample.
When combined with standard loss functions, in particular DSC, mL1-ACE demonstrates superior calibration performance on the BraTS2021 dataset [1], significantly surpassing conventional post-hoc calibration methods like Ts. Additionally, we introduce dataset reliability histograms, offering a comprehensive visual tool for evaluating and communicating model calibration across image segmentation datasets, enhancing transparency in reporting model reliability beyond predictive performance.
@@ 2 Method
where $n^{c}_{m}$ is the number of in bin $B^{c}_{m}$, the $m^{\textrm{th}}$ bin for class $c$, and $N$ is the total number of voxels for an image. More formally, we define the observed foreground frequency $o^{c}_{m}$ and the expected foreground probability $e^{c}_{m}$ as:
@@ 4 Results
Figure 3: Comparative analysis of DSC against calibration metrics for average tumour component of BraTS 2021 dataset. Statistically significant difference (p < 0.01) for all metrics.
@@
Loss Function ET TC WT Avg
CE 0.82 ± 0.19 0.87 ± 0.18 0.90 ± 0.11 0.86 ± 0.14
CE + L1-ACE 0.80 ± 0.20 0.86 ± 0.19 0.87 ± 0.12 0.84 ± 0.15
DSC 0.84 ± 0.17 0.87 ± 0.20 0.89 ± 0.12 0.87 ± 0.14
DSC + L1-ACE 0.83 ± 0.18 0.88 ± 0.18 0.89 ± 0.11 0.87 ± 0.13
DSC-CE 0.85 ± 0.17 0.89 ± 0.17 0.90 ± 0.11 0.88 ± 0.13
DSC-CE + L1-ACE 0.82 ± 0.20 0.86 ± 0.21 0.88 ± 0.13 0.85 ± 0.16
We show the effect of Ts and mL1-ACE on the calibration of the models in Table 2. Our auxiliary loss has little effect on the calibration of the model trained with CE. Our loss shows the biggest improvement in calibration when paired with the DSC loss function, with a 45% reduction in ACE (p < 0.01). Overall the use of Ts shows marginal improvement in calibration, with the exception of CE loss where it has a small detrimental effect. Similar conclusion can be drawn for MCE and ECE metrics, shown in Table S1 and Table S2 in the supplementary material. We also note that the number of bins used to calculate our calibration metrics, has little effect, with ACE increasing by less than 3% when increasing the bins from 10 to 100 (p < 0.01), as shown in supplementary Fig. S1.
@@ 5.0.1 Acknowledgements
[…]
TV is co-founder and shareholder of Hypervision Surgical. BG is a part-time employee of Kheiron Medical Technologies and HeartFlow with stock options as part of the standard compensation package. The authors have no other competing interests to declare that are relevant to the content of this article.
Supplementary Material
Table S1: Maximum Calibration Error
ET TC WT Avg
CE 0.26 ± 0.19 0.28 ± 0.18 0.23 ± 0.12 0.25 ± 0.12
CE + Ts 0.30 ± 0.16 0.29 ± 0.16 0.25 ± 0.10 0.28 ± 0.10
CE + L1-ACE 0.24 ± 0.19 0.26 ± 0.18 0.20 ± 0.15 0.23 ± 0.14
CE + L1-ACE + Ts 0.23 ± 0.15 0.25 ± 0.16 0.20 ± 0.12 0.23 ± 0.11
Dice 0.59 ± 0.15 0.61 ± 0.15 0.55 ± 0.13 0.58 ± 0.11
Dice + Ts 0.56 ± 0.15 0.58 ± 0.15 0.51 ± 0.14 0.55 ± 0.11
Dice + L1-ACE 0.29 ± 0.18 0.30 ± 0.17 0.23 ± 0.14 0.27 ± 0.13
Dice + L1-ACE + Ts 0.23 ± 0.17 0.25 ± 0.16 0.21 ± 0.11 0.23 ± 0.12
Dice-CE 0.46 ± 0.17 0.46 ± 0.17 0.36 ± 0.17 0.43 ± 0.13
Dice-CE + Ts 0.26 ± 0.19 0.32 ± 0.19 0.19 ± 0.12 0.26 ± 0.13
Dice-CE + L1-ACE 0.29 ± 0.21 0.30 ± 0.20 0.21 ± 0.14 0.27 ± 0.15
Dice-CE + L1-ACE + Ts 0.25 ± 0.19 0.27 ± 0.18 0.20 ± 0.11 0.24 ± 0.13
Table S2: Expected Calibration Error (×$\times 10^{-4}$)
ET TC WT Avg
CE 4.21 ± 7.49 5.66 ± 10.5 9.71 ± 11.3 6.53 ± 7.53
CE + Ts 7.77 ± 7.16 10.2 ± 10.0 19.1 ± 9.45 12.4 ± 6.93
CE + L1-ACE 4.00 ± 4.75 6.16 ± 9.41 12.0 ± 12.9 7.38 ± 6.94
CE + L1-ACE + Ts 21.1 ± 4.69 27.2 ± 6.65 62.0 ± 11.1 36.8 ± 5.69
Dice 5.09 ± 7.67 7.48 ± 16.5 15.5 ± 17.8 9.34 ± 11.1
Dice + Ts 4.95 ± 7.63 7.36 ± 16.5 15.0 ± 17.8 9.10 ± 11.1
Dice + L1-ACE 3.47 ± 7.56 4.50 ± 8.45 10.1 ± 12.8 6.03 ± 6.84
Dice + L1-ACE + Ts 3.27 ± 7.17 4.41 ± 8.00 12.0 ± 10.9 6.54 ± 5.84
Dice-CE 4.40 ± 7.57 5.70 ± 12.2 11.6 ± 14.4 7.21 ± 8.62
Dice-CE + Ts 3.84 ± 7.43 5.20 ± 11.8 10.2 ± 13.4 6.41 ± 8.11
Dice-CE + L1-ACE 3.91 ± 8.09 6.04 ± 14.6 10.3 ± 14.8 6.75 ± 9.84
Dice-CE + L1-ACE + Ts 4.23 ± 7.68 6.56 ± 14.0 13.9 ± 12.6 8.23 ± 8.83
@@
Figure S1: Comparison of calibration metrics using different number of bins when evaluated on the DSC + mL1-ACE loss model. Both ACE and ECE are very stable with respect to number of bins (5, 10, 20, 50, 100, 1000). MCE shows a larger increase from 100 to 1000 bins, this is due to the presence of empty bins, a likely occurrence when using such a high number of bins.